OpenStudy (anonymous):

graph the function f(x)=-4x

OpenStudy (anonymous):

This is tough to type so you'll have to do the work: First substitute a few values for the x, say -2, -1, 0, 1 and 2 You'll get 4 values for f(x) -4(-2)=8 -4(-1)=4 -4(0)=0 -4(1)=-4 -4(2)=-8 Next in a graph you plot those 5 points (-2,8), (-1,4), (0,0), (1,-4) and (2,-8) then just draw a line through all those points and you're set. If the line is a little short for your chart, just use a ruler and extend it as far as you like

OpenStudy (dumbcow):

good answer, if you know about slopes yet you could also start at origin and count down 4 over 1